Circular No. 6138

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

SUPERNOVA 1995F IN NGC 2726 D. J. Lane and P. Gray, Burke-Gaffney Observatory, Saint Mary’s

University, report their discovery via CCD on Feb. 10–11 of a supernova located ∼ 2′′ east and 1′′ south of the center of NGC 2726 (α = 9^{h}04^{m}.9, δ = +59^{o}56′, equinox 2000.0). An unfiltered frame on Feb. 14.042 UT showed the object at mag 14.7. A foreground star of mag 14.9 is located 19′′.6 west and 14′′.9 south of SN 1995F. A. V. Filippenko and A. J. Barth, University of California at Berkeley, report that preliminary inspection of uncalibrated CCD spectra (range 310–1050 nm, resolution 0.7–1.5 nm) obtained on Feb. 24 with the 3-m Shane reflector at Lick Observatory confirms that this object is a supernova, most likely of type Ic (but possibly Ib), roughly 2–3 weeks past maximum brightness.

SUPERNOVA 1995G IN NGC 1643 Robert O. Evans, Coonabarabran, N.S.W., communicates his visual

discovery on Feb. 23.5 UT with J. Shobbrook and S. Beaman, using the Australian National University’s 1-m reflector at Siding Spring, of an apparent supernova (mag 15.5) in NGC 1643 (α = 4^{h}43^{m}.6, δ = −5^{o}19′, equinox 2000.0). SN 1995G is located ∼ 3′′ east and 15′′ north of the galaxy’s center. A confirming 4415 film was exposed for 10 min by P. Cass with the U.K. Schmidt Telescope on Feb. 24.45, aided by comparison with an older red survey plate.

SUPERNOVA 1995E IN NGC 2441 D. D. Balam relays the following precise position for SN 1995E from

and observation made on Feb. 23.267 UT with the 1.82-m Plaskett telescope at the Dominion Astrophysical Observatory, Victoria: α = 7^{h}51^{m}56^{s}.75, δ = +73^{o}00′34′′.6 (equinox 2000.0); the corresponding offset from the galaxy’s nucleus is 8′′.6 east and 20′′.6 south. Preliminary photometry from a leastsquares fit of 11 GSC stars: V = 15.9 ± 0.1.

Corrigendum. On IAUC 6137, line 2, for Gabrielcic read Gabrijelcic

SUPERNOVA 1995C IN ANONYMOUS GALAXY Filippenko and Barth report that inspection of a CCD spectrum (ob-

tained as above; range 420–700 nm, resolution 0.7 nm) with the 3-m Shane reflector at Lick Observatory reveals that SN 1995C (cf. IAUC 6133) is of type Ia, ∼ 3 weeks past maximum brightness.

1995 February 24 Daniel W. E. Green